OpenStudy (anonymous):

PLEASE HELP!!!! How do i turn: f(x)=28.65e^(-64x) into a logarithmic function???

OpenStudy (amistre64):

this one?

OpenStudy (amistre64):

\[y = 28.65 e^{-64x}\] ln it all \[ln(y) = ln(28.65 e^{-64x})\] \[ln(y) = ln(28.65)+ln(e^{-64x})\] \[ln(y) = ln(28.65)-64x\] \[ln(y) -ln(28.65)=-64x\] \[ln(\frac{y}{28.65})=-64x\] \[-\frac{1}{64}ln(\frac{y}{28.65})=x\] is the only thing I can think of for it

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (amistre64):

yep, hope it helps :)